I feel a little confused about this problem
x/6 + 3 =9. Am I supposed to divide the six into the x and then add the 3? The answer I was getting was 3, and then I found out that it was wrong and I would like to know where I went wrong.
-----------------------------------------------------------------
An arithmitic student needs an average of 70 to recieve credit. She scored 82,64 and 98 on the first three exams. I am looking for a equality representing the score she must get on the last exam to recieve credit for the course.
--------------------------------------------------------------------
Is 15 a solution to the equation 8-x=7?
--------------------------------------------------------------------
Solve for x, 3x-2y=6.
I forgot how to solve a problem like this one. I just need a refresher so I can get the other ones solved.

Answer by stanbon(75887)

x/6 + 3 =9. Am I supposed to divide the six into the x and then add the 3? The answer I was getting was 3, and then I found out that it was wrong and I would like to know where I went wrong.
When you have an equation multiply thru by the least common multiple of the
denominators to cancel all the denominators
Your Problem:
Multiply thru by 6 to get:
x+18 = 54
x= 36
-----------------------------------------------------------------
An arithmitic student needs an average of 70 to recieve credit. She scored 82,64 and 98 on the first three exams. I am looking for a equality representing the score she must get on the last exam to recieve credit for the course.
Let "x" be the score she needs.
-Average is the sum of the scores divided by the number of scores:
EQUATION:
(82+64+98+x)/4 = 70
244+x = 280
x = 36 (She needs to get at least 36 on the 4th test)
===========
--------------------------------------------------------------------
Is 15 a solution to the equation 8-x=7?
Substitute and see if you get a true statement:
8-15 = 7
-7=7
15 is not a solution
--------------------------------------------------------------------
Solve for x, 3x-2y=6.
3x = 2y+6
x= (2/3)x + 3
========================
Cheers,
Stan H.